OpenStudy (anonymous):

(5 cot x + 1)(sin x + 7) = 0 approximate nearest degree within [0, 360)

OpenStudy (cwrw238):

5 cot x + 1 = 0 cot x = -0.2 tan x = -5

OpenStudy (cwrw238):

sin x + 7 = 0 gives no real values of x

OpenStudy (cwrw238):

tan x = -5 gives 2 values in 2nd and 4th quadrant x = 180 - 78.69 = 101.31 or 360 - 78.69 = 281.31 to nearest degree solution is 101 , 281 degrees